A coin is tossed 200 times and it lands heads up 115 times. At the 5% significance level, is there sufficient evidence to claim that the coin is not equally likely to land heads or tails?

(a) State the null & alternative hypotheses.
H0:
Ha:
p-value = _______

Decision: ___________ (Reject, Fail to reject) H0.

Interpretation: In Words



Answer by stanbon(75887)

A coin is tossed 200 times and it lands heads up 115 times. At the 5% significance level, is there sufficient evidence to claim that the coin is not equally likely to land heads or tails?
(a) State the null & alternative hypotheses.
H0: p = 1/2
Ha: p # 1/2 (claim)
-------------------
p-hat = 115/200 = 0.575
-----
z(0.575) = (0.575-0.5)/sqrt(0.5^2/200) = 2.1213
---
p-value = 2*p(z > 2.1213) = 2*normalcdf(2.1213,100) = 0.0339
------
Decision: Since the p-value is less than 5%, Reject H0.
----
Interpretation: In Words
Decision:: The test results support the claim that the
coin is not equally likely to land heads or tails at
the 5% significance level.
---------------
Cheers,
Stan H.
---------